Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0. The solutions are -6,3.
Found 2 solutions by mananth, algebrahouse.com:
Answer by mananth(16946)

solution (-6,3)
equation
(x+6)(x-3)=0
x^2+3x-18=0

Answer by algebrahouse.com(1659)

Start with the solutions and work backwards to the equation.

x = -6 or x = 3
x + 6 = 0 or x - 3 = 0 {added 6 and subtracted 3, respectively}
(x + 6)(x - 3) = 0 {set the two binomials, multiplied together and equal to 0}
x² + 3x - 18 = 0 {multiplied with foil method}

For more help from me, visit: www.algebrahouse.com